Circular No. 8407

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 R3 (LINEAR-NEAT) On Sept. 15 the NEAT project at Jet Propulsion Laboratory reported

to the Minor Planet Center observations from Sept. 13 of a cometary object observed at Palomar in collaboration with M. Brown, California Institute of Technology. Observations with the 1.2-m Schmidt telescope showed the object to be diffuse with a coma ∼ 5′′ across and condensation of 2′′ but no evident tail. The Minor Planet Center then established that the object was identical with an unremarkable, presumed main-belt minor planet for which it had linked observations from Sept. 10 and 13 reported by LINEAR. The two “discovery” observations are as follows:

2004 UT α_{2000} δ_{2000} Mag. Observer
Sept.10.29965 0^{h}37^{m}23.^{s}60 +14^{o}18′11′′.9 20.1 LINEAR
13.26186 0 35 27.70 +14 21 11.6 19.1 NEAT

Following placement on “The NEO Confirmation Page” further observations were received, among them a report from G. Jones, Tucson, AZ (0.32-m Ritchey-Chretien reflector), to the effect that on Sept. 16.20 UT the object had a coma of diameter 3′′ and a 10′′ tail in p.a. 230^{o}, as well as one from J. Young, Table Mountain Observatory (0.6-m reflector), remarking that on Sept. 16.23 there was a 5′′ coma (with no apparent central condensation) and a fan-shaped tail between p.a. 210^{o} and 270^{o} extending ∼ 8′′; a straight, narrow spike extended 40′′ in p.a. 236^{o}.

The full set of 32 astrometric observations, equinox-2000.0 orbital elements (T = 2004 May 23, ω = 5^{o}, Ω = 319^{o}, i = 8^{o}, q = 2.14 AU, P = 7.6 years), and an ephemeris are given on MPEC 2004-S02.

SUPERNOVA 2004eq IN ESO 404-12 K. Huang, Y. Chen, Z. Lin, H. Lin, M. Tsai, and W. Ip, Institute of

Astronomy, National Central University, Taiwan; and Y. Qiu and J. Hu, National Astronomical Observatories of China, report the discovery of an apparent supernova (mag ∼ 16.8) on unfiltered CCD images taken on July 30.72 UT with the 1-m Lulin Observatory telescope. SN 2004eq is located at α = 21^{h}27^{m}35^{s}.43, δ = −34^{o}34′58′′.0 (equinox 2000.0), which is 28′′.7 east and 5′′.0 north of the center of ESO 404-12. Nothing is visible at this position on images taken on July 10.667. Additional approximate unfiltered magnitudes for the new object: July 31.68, 16.2; Aug. 13.62, 13.4; Sept. 1.56, 14.2; 5.70, 14.3; 19.65, 15.3; 20.59, 15.3.

2004 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green